# Influence of intraspecies interactions on the diversity of the wetting phase diagram in dilute ternary Bose-Einstein condensates

### v1

@@ 1 Introduction
The wetting phenomenon in Bose-Einstein condensates (BECs) was first predicted theoretically in 2004 [8], in the context of a two-component BEC confined by an idealized hard-wall (optical wall). Despite the theoretical interest it has garnered, this phenomenon has yet to be observed experimentally, even after more than two decades. Several hypotheses have been advanced to explain this absence, accompanied by a range of theoretical and experimental efforts aimed at overcoming the associated challenges. One key limitation is the ideal nature of the hard-wall potential, which cannot be realized in practical experimental setups. To address this issue, alternative approaches have been proposed, including the replacement of the hard-wall confinement with a more experimentally accessible soft-wall potential [9], and investigations into the thermodynamic behavior of the prewetting transition [10].
A more recent development involves the substitution of the conventional hard- or soft-wall boundary by a third constituent within a ternary (three-component) BECs system [11]. This configuration is characterized not only by the atomic masses $m_{i},(i=1,2,3)$, but also by six intrinsic interaction parameters: three intraspecies scattering lengths $a_{ii}$, and three interspecies scattering lengths $a_{ij}=a_{ji},(i,j=1,2,3,i\neq j)$. Additionally, the system comprises three independent particle densities $n_{i}$. The presence of a third component in BEC systems markedly increases the number of experimentally controllable parameters, particularly the relative coupling constants $K_{13},K_{23}$ and the healing length ratios $\xi_{i}/\xi_{j}$ (see below). These parameters can be tuned either independently or in conjunction, in BEC mixtures composed of two hyperfine states of a single atomic species or of two distinct atomic species [12, 13, 14, 15, 16]. A notable example involves laser-induced shifts of magnetic Feshbach resonances in 87Rb, which permit precise tuning of the scattering lengths while maintaining significantly lower particle loss rates than those associated with optical Feshbach resonances [14]. Furthermore, multicomponent systems featuring both ferromagnetic and antiferromagnetic spin–spin interactions offer an extended degree of control [16]. In mixtures of 87Rb and 133Cs, the interspecies scattering length—which characterizes atomic interactions—has been successfully manipulated using magnetic Feshbach resonances in magnetic quadrupole traps, followed by confinement in levitated optical traps [12], or through the use of cigar-shaped trapping geometries [13]. This enhanced tunability gives rise to increasingly rich and intricate wetting phase diagrams, underscoring the complexity and flexibility of multicomponent BEC systems.
@@ 2 Wave functions of a dilute ternary BEC in double parabola approximation
Here $\hbar$ denotes the reduced Planck constant, $\mu_{i}$ is the chemical potential of the $i$-th component, and $\psi_{i}(\vec{r})$ is the macroscopic wave function associated with that component. In the absence of superfluid flow, the wave functions can be taken as real and serve as order parameters. The local condensate density of each component is then defined by $n_{i}(\vec{r})\equiv|\psi_{i}(\vec{r})|^{2}$. The interaction strengths are characterized by the coupling constants. In the dilute BECs, where the $s$-wave scattering lengths $a_{ii}$ and $a_{ij}$ are much smaller than the average interatomic spacing, the interatomic interactions can be modeled as contact potentials. In this limit, the intraspecies coupling constants take the form
@@
Here we use $\widetilde{V}=V/(g_{11}n_{1}^{2}/2)$, which is dimensionless potential. From (12) the coupled GP equations in (2) become
@@
- $(\frac{\mu_{3}}{\bar{\mu}_{3}}K_{13}-1)\tilde{\psi}_{1}^{2}-(\frac{\mu_{3}}{\bar{\mu}_{3}}K_{23}-1)\tilde{\psi}_{2}^{2}.$
@@
$\tilde{\psi}_{1}$ $=$ $2B_{1}\,\sinh(\sqrt{\frac{\mu_{3}}{\bar{\mu}_{3}}K_{13}-1}\,\frac{\xi_{2}}{\xi_{1}}\tilde{z}),\;\mbox{for}\;\tilde{z}<0,$
$\tilde{\psi}_{2}$ $=$ $-2C_{2}\,\sinh(\sqrt{\frac{\mu_{3}}{\bar{\mu}_{3}}K_{23}-1}\,\tilde{z}),\;\mbox{for}\;\tilde{z}>0,$
$\tilde{\psi}_{3}$ $=$ $\sqrt{\frac{\mu_{3}}{\bar{\mu}_{3}}}+B_{3}\,\exp{(\sqrt{2}\sqrt{\frac{\mu_{3}}{\bar{\mu}_{3}}}\frac{\xi_{2}}{\bar{\xi}_{3}}\tilde{z})}+C_{3}\,\exp{(-\sqrt{2}\sqrt{\frac{\mu_{3}}{\bar{\mu}_{3}}}\frac{\xi_{2}}{\bar{\xi}_{3}}\tilde{z})}$. (18)
@@ 3.1 Wetting phase diagram
The interfacial tension at the interface 1-2 with presence of the surfactant is defined as
@@ 4 Conclusions
In this work, we have investigated the wetting phase transition in a dilute ternary Bose–Einstein condensate (BEC). Motivated by experimental feasibility, we focused on scenarios in which the intraspecies atomic interactions are tunable, while the interspecies interactions are held fixed.
@@
    The locus of the degenerate points in plane of the relative coupling constants has been investigated.

### v2

@@ 1 Introduction
The wetting phenomenon in Bose-Einstein condensates (BECs) was first predicted theoretically in 2004 [8], in the context of a two-component BECs confined by an idealized hard-wall (optical wall). Despite the theoretical interest it has garnered, this phenomenon has not yet to be observed experimentally, even after more than two decades. Several hypotheses have been advanced to explain this absence, accompanied by a range of theoretical and experimental efforts aimed at overcoming the associated challenges. One key limitation is the ideal nature of the hard-wall potential, which cannot be realized in practical experimental setups. To address this issue, alternative approaches have been proposed, including the replacement of the hard-wall confinement with a more experimentally accessible soft-wall potential [9], and investigations into the thermodynamic behavior of the prewetting transition [10].
A more recent development involves the substitution of the conventional hard- or soft-wall boundary by a third constituent within a ternary (three-component) BECs system [11]. This configuration is characterized not only by the atomic masses $m_{i},(i=1,2,3)$, but also by six intrinsic interaction parameters: three intraspecies scattering lengths $a_{ii}$, and three interspecies scattering lengths $a_{ij}=a_{ji},(i,j=1,2,3,i\neq j)$. Additionally, the system comprises three independent particle densities $n_{i}$. The presence of a third component in BEC systems markedly increases the number of experimentally controllable parameters, particularly the relative coupling constants $K_{13},K_{23}$ and the healing length ratios $\xi_{i}/\xi_{j}$ (see below). These parameters can be tuned either independently or in conjunction, in BECs mixtures composed of two hyperfine states of a single atomic species or of two distinct atomic species [12, 13, 14, 15, 16]. A notable example involves laser-induced shifts of magnetic Feshbach resonances in 87Rb, which permit precise tuning of the scattering lengths while maintaining significantly lower particle loss rates than those associated with optical Feshbach resonances [14]. Furthermore, multicomponent systems featuring both ferromagnetic and antiferromagnetic spin–spin interactions offer an extended degree of control [16]. In mixtures of 87Rb and 133Cs, the interspecies scattering length, which characterizes atomic interactions, has been successfully manipulated using magnetic Feshbach resonances in magnetic quadrupole traps, followed by confinement in levitated optical traps [12], or through the use of cigar-shaped trapping geometries [13]. This enhanced tunability gives rise to increasingly rich and intricate wetting phase diagrams, underscoring the complexity and flexibility of multicomponent BEC systems.
@@ 2 Wave functions of a dilute ternary BEC in double parabola approximation
Here $\hbar$ denotes the reduced Planck constant, $\mu_{i}$ is the chemical potential of the $i$-th component, and $\psi_{i}(\vec{r})$ is the macroscopic wave function associated with that component. In the absence of particle flow, the wave functions can be taken as real and serve as order parameters. The local condensate density of each component is then defined by $n_{i}(\vec{r})\equiv|\psi_{i}(\vec{r})|^{2}$. The interaction strengths are characterized by the coupling constants. In the dilute BECs, where the $s$-wave scattering lengths $a_{ii}$ and $a_{ij}$ are much smaller than the average interatomic spacing, the interatomic interactions can be modeled as contact potentials. In this limit, the intraspecies coupling constants take the form
@@
Here we use $\widetilde{V}\equiv V/(g_{11}n_{1}^{2}/2)$, which is dimensionless potential. From (12) the coupled GP equations in (2) become
@@
- $\left(\frac{\mu_{3}}{\bar{\mu}_{3}}K_{13}-1\right)\tilde{\psi}_{1}^{2}-\left(\frac{\mu_{3}}{\bar{\mu}_{3}}K_{23}-1\right)\tilde{\psi}_{2}^{2}.$
@@
$\tilde{\psi}_{1}$ $=$ $2B_{1}\,\sinh\left(\sqrt{\frac{\mu_{3}}{\bar{\mu}_{3}}K_{13}-1}\,\frac{\xi_{2}}{\xi_{1}}\tilde{z}\right),\;\mbox{for}\;\tilde{z}<0,$
$\tilde{\psi}_{2}$ $=$ $-2C_{2}\,\sinh\left(\sqrt{\frac{\mu_{3}}{\bar{\mu}_{3}}K_{23}-1}\,\tilde{z}\right),\;\mbox{for}\;\tilde{z}>0,$
$\tilde{\psi}_{3}$ $=$ $\sqrt{\frac{\mu_{3}}{\bar{\mu}_{3}}}+B_{3}\,\exp{\left(\sqrt{\frac{2\mu_{3}}{\bar{\mu}_{3}}}\frac{\xi_{2}}{\bar{\xi}_{3}}\tilde{z}\right)}+C_{3}\,\exp{\left(-\sqrt{\frac{2\mu_{3}}{\bar{\mu}_{3}}}\frac{\xi_{2}}{\bar{\xi}_{3}}\tilde{z}\right)}$. (18)
@@ 3.1 Wetting phase diagram
The interfacial tension at the 1-2 interface with presence of the surfactant is defined as
@@ 4 Conclusions
In this work, we have investigated the wetting phase transition in a dilute ternary BECs. Motivated by experimental feasibility, we focused on scenarios in which the intraspecies atomic interactions are tunable, while the interspecies interactions are held fixed.